Solve

x ^2y'′ + 2xy'− 12y = x^3log x.



1
Expert's answer
2022-03-16T14:33:12-0400

"a_0\lambda(\lambda-1)(\lambda-2)...(\lambda-n+1)+...a_{n-2}\lambda(\lambda-1)+a_{n-1}\lambda+a_n=0"

x=eu

"(\lambda-1)\lambda+2\lambda-12=0"

"\lambda^2+\lambda-12=0"

"\lambda_1=3"

"\lambda_2=-4"

"y''+y'-12y=ue^{3u}"

"y_0=Ce^{3u}+\frac{C_1}{e^{4u}}"

Solution for "ue^{3u}"

a+bi=3, then s=1

"y_1=u(Au+B)e^{3u}"

"y'_1=(3Au^2+(3B+2A)u+B)e^{3u}"

"y''_1=(9Au^2+(9B+12A)u+6B+2A)e^{3u}"

"14Aue^{3u}+(7B+2A)e^{3u}=ue^{3u}"


A=1/14

B=-1/49

"y_1=(\frac{u}{14}-\frac{1}{49})ue^{3u}"

"y=(\frac{u}{14}-\frac{1}{49})ue^{3u}+Ce^{3u}+\frac{C_1}{e^4u}"